OpenStudy (anonymous):

What is the 32nd term of the arithmetic sequence where a1 = 4 and a6 = 24 ?

OpenStudy (anonymous):

a=4 a6=a+(6-1)d=24 4+5d=24 5d=20 d=4... a32=a+31d=4+31*4=128

OpenStudy (anonymous):

thank youuu!